OpenStudy (anonymous):

An equation is shown below: 5(2x - 3) = 5 Part A: How many solutions does this equation have? Part B: What are the solutions to this equation? Show your work.

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

$$\Huge a( b + c ) = ab +ac$$

OpenStudy (skullpatrol):

$$\Huge a(b-c)=ab-ac$$

OpenStudy (skullpatrol):